OpenStudy (anonymous):

x + 4 in parentheses over 2 = 7 (Solve) A. x = 3 B. x = 5 C. x = 6 D. x = 10

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

how are you sure?

OpenStudy (mokeira):

\[\frac{ (x+4) }{ 2 }=7\] multiply both sides by 2 \[x+4=14\] x=14-4 x=10

OpenStudy (anonymous):

because its just right

OpenStudy (anonymous):

I did the problem over and i did get x= 10 Thank you syriafr for the effort! and thank you mokeira swell :)

OpenStudy (mokeira):

@syriafr it cant just be right explain to @ConnorRocheo how you did it..understanding how to do it is more important than the answer

OpenStudy (mokeira):

@ConnorRocheo so you were right! Nice!! welcome :)

OpenStudy (anonymous):

@Mokeira can you help me with some other problems?

OpenStudy (mokeira):

i can try